OpenStudy (anonymous):

i need help with this question.................. The slope of the line whose equation is x + 2y = 3 is -1/2 -1 -2

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

Make the equation look like this

OpenStudy (anonymous):

do u understand?

OpenStudy (anonymous):

the slope is -1/2

OpenStudy (anonymous):

2y=3-x 2y=-x+3 y=-(1/2)X+(3/2)

OpenStudy (anonymous):

m=-1/2

OpenStudy (anonymous):

Determine the slope of the given line. 2y = x-3 y = (1/2)x - (3/2) slope of the given line is 1/2 Therefore slope of any perpendicular line must be -2 Form of needed equation: y = mx + b y = -7 when x = 4 and the slope is -2 Solve for "b": -7 = -2*4 + b b = 1 Equation: y = -2x +1

OpenStudy (anonymous):

so , it will be m=-1/2

OpenStudy (anonymous):

@ryanvarghese12790 the given equation is not 2y=x-3, it is 2y = -X + 3

OpenStudy (anonymous):

i don't know how y came up with -1/2

OpenStudy (anonymous):

May I help you @khofer ?

OpenStudy (anonymous):

Look : If we want to have a standard form should be have y alone . So have this : y=mx+d So y is alone means nothing is before or after that ... So we have : x + 2y = 3 2y=-x+3 But 2 is extra .. So delete it from this way ...:(Divided on 2) y=-x/2+3/2 Got it ?:) @khofer

OpenStudy (anonymous):

yes thank u....

OpenStudy (anonymous):

You re welcome ! I wish my help can be use full caz last answers was clear and good .